A hotel offers two activity packages. One costs $192 and includes 3 hours of horseback riding and 2 hours of parasailing. The second costs $213 and includes 2 hours of horseback riding and 3 hours of parasailing. What is the cost for 1 hour of each activity?”

1
Expert's answer
2021-02-24T06:48:50-0500

If the cost of horseback riding is h and the cost of parasailing is p, then we get a system of the two equations:

3h + 2p = 192,

2h + 3p = 213.

From the first equation p = 96 - 1.5h. If we put it into the second equation, then:

2h + 288 - 4.5h = 213,

2.5h = 75,

h = 30,

p = 96 - 1.5×30 = 51.